Central Park, which lies in the middle of Manhatten, is rectangular shaped and has a 6 mile perimeter. The length is 5 times the width. What are the dimensions of the park?
Answer by Cromlix(4381)

Perimeter = 2 x length + 2 x width
Width = x
Length = 5x
6 miles = 2(5x) + 2(x)
6 miles = 10x + 2x
6 miles = 12x
x = 6/12
x = 1/2 mile
Width = 1/2 mile
Length = 5/2 miles or 2.5 miles
Perimeter = 2 x length + 2 x width
= 2 x 2.5 + 2 x 0.5
= 5 miles + 1 mile
= 6 miles
Hope this helps.
:-)